I follow along to a point, but some things stump me. Would you show me how to solve lnA = lnH-ln(1-H)
Answer by scott8148(5880)

taking antilog __ A=H/(1-H)

multiplying by 1-H __ A-AH=H

adding AH __ A=H+AH

factoring __ A=H(1+A)

dividing (1+A) __ A/(1+A)=H